A vector A  has components Ax = 12.0 m and Ay= 5.00 m. What is the angle that vector A  makes with the x-axis?


1
Expert's answer
2021-09-28T17:11:21-0400

"tan \theta= Ay/Ax=5/12=0.4167"

"\theta=tan^{-1}0.4167=22.62^o"